Grade 12Physical Chemistry

For first order reaction the rate of reaction at 10 minute is 0.04/sec and at 20minute it is 0.03/sec.Determine the half life period of this reaction

harshili
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
Dear student
 
 
U have to just put the values into the formulae for the first order eqn. that is:
K = (2.303/t2 –  t1)log(Q1/Q2)
   =(2.303/(10)llog(0.04/0.03)
After solving,
K = 0.02876 /min
thalf = 0.693/k = > 24 min approx.